OpenStudy (anonymous):

given f(x) = two x minus two over four solve for f^1(3)

OpenStudy (anonymous):

same as solving \(\frac{2x-2}{4}=7\) but instead solve \[\frac{2y-2}{4}=x\] for \(x\)